# The JWST Galactic Center Survey A White Paper

### v6

@@ 2.2 The unique observational capabilities of JWST in Galactic Center context
Angular resolution: Crowding is the main factor that limits the detection of faint stars at the GC, such as old stars near the main sequence turn-off and low-mass YSOs (Fig. 1). JWST provides a ten times higher angular resolution than Spitzer and can deliver almost the resolution of adaptive optics-assisted cameras at ground-based 8 m-class telescopes. However, since precision measurements with the latter are strongly limited by anisoplanatic effects, their useful field-of-view is only approximately 0.25 arcmin, while the effective field-of-view of NIRCam is roughly 40 times larger in each channel.
@@
Ground-based, adaptive optics-assisted instruments on 10m-class telescopes are no alternative to JWST, because (1) they can cover at most arcminute-wide fields with a systematically changing PSF due to anisoplanatic effects, (2) zero point variations, mostly due to PSF instability, are a problem for multi-epoch observations, and (3) they are about times less sensitive at m than JWST. (1) implies that accurate photometry and astrometry can only be performed over very small fields (0.25 arcmin), thus making any observations of large fields highly time-consuming. (3) means that less than 10% of the stars detected at m can be detected at m from the ground.
@@ 2.3 JWST will produce transformational measurements
With its high sensitivity JWST can potentially detect YSOs down to L 0.1 in fields where detection completeness is not limited by crowding, such as star-forming regions associated with dark clouds. Current data, mainly from the Spitzer Space Telescope, limit detections to massive YSOs, which are rare and, in the Spitzer data, confused with other sources.
@@
With a minimum of two NIRCam epochs, we can measure the proper motions of an estimated 10 million stars at the GC. No existing or planned mission (including Gaia NIR) is or will be able to achieve this. A five year time base will allow proper motion measurements with an accuracy of at least mas yr (10 km s). Here we have considered mag stars and have assumed that alignment between two observing epochs can be done with an accuracy of 1 mas. The proper motions will link the Galactic Center to the rest of the Milky Way where proper motions are measured with Gaia and the Vera Rubin observatory. Proper motions are key to disentangling different stellar structures (bar, nuclear stellar disk, nuclear star cluster), identifying accretion events onto the nuclear star cluster and understanding the structure of the nuclear stellar disk (is it an inner bar?). Even a single epoch JWST survey will already yield about one million proper motions by combination with archival HST (e.g. Dong et al. 2011; Libralato et al. 2021) and ground-based images (from the GALACTICNUCLEUS survey).
@@ 2.4 Survey requirements to address the key open questions
Filters: F140M, F187N, F210M, F405N and F480M to enable accurate dereddening, minimize saturation, trace ISM features, and obtain the best possible color-magnitude diagrams (CMDs). Since confusion is less of an issue at the shortest wavelengths, where extinction is significantly higher, an alternative to F140M imaging with NIRCam may be F127M imaging with WFC3/HST.
@@ 2.5 Synergy with other facilities
ALMA has recently completed the ALMA CMZ Exploration Survey (ACES) in the 3.2 mm continuum and over a dozen spectral lines with 2 angular resolution.
@@
The Roman Space Telescope (launch 2027) offers a compelling opportunity to perform a high-cadence survey of the GC region at HST-like resolution (e.g. Terry et al. 2023) for time-domain science, albeit with a significantly worse angular resolution than JWST (thus hitting the crowding limit at brighter magnitudes). The proposed JWST observations will complement such a survey by providing star-by-star extinction measurements, much-improved depth in highly crowded and/or extinguished regions, identification of stars whose photometry/astrometry is significantly biased by nearby companions, and an epoch of high-precision astrometry for proper motion measurements.
@@
X-ray observatories have revealed the presence of an outflow from the Galactic Center (e.g. Ponti et al. 2019). The proposed observations may allow us to connect star formation regions to outflows and thus support the interpretation of X-ray observations.
@@ 3 Solving multiple key science problems with a single data set
The 50050 pc CMZ contains 2 - 7 M of dense molecular gas, corresponding to % of our Galaxy’s molecular gas. Most of the mass is distributed in a roughly ring-like/elliptical shape of 100 pc () radius around Sgr A*, the so-called 100 pc-stream (Fig. 2, e.g. Henshaw et al. 2022). The CMZ includes exceptional ISM features such as the M “Brick” (Longmore et al. 2012) and the M Sgr B2 cloud (de Pree et al. 1995; Schmiedeke et al. 2016; Ginsburg et al. 2018, and references therein). The CMZ occupies of the volume of the Milky Way’s disk and reaches a mean molecular gas density times higher than in the Galactic disk.
The NSD consists of M of stars in a flat, rotating, disk-like structure that has an effective radius of 100 pc and a scale-height of 40 pc (Launhardt et al. 2002; Schönrich et al. 2015; Henshaw et al. 2022; Sormani et al. 2022). With a mass of M and an effective radius of about 4 pc, the NSC is the densest concentration of stars in the Galaxy (Schödel et al. 2014b; Feldmeier-Krause et al. 2017; Neumayer et al. 2020). The majority of the stars in the NSD and NSC, , appear to have formed more than 8-10 Gyr ago (Nogueras-Lara et al. 2020a; Schödel et al. 2020) and are the most metal-rich population of old stars in the Galaxy (Schultheis et al. 2021). At the kinematic and density center of the NSC lies the M super-massive black hole Sagittarius A* (Sgr A*) (Ghez et al. 2008; Gravity Collaboration et al. 2020).
Averaged by volume, the mean star formation rate at the GC is one to two orders of magnitude higher than in the Galactic disk, with extreme conditions that resemble those in high redshift star-forming galaxies (Kruijssen & Longmore 2013). Ongoing and intense past star formation activity is witnessed by a broad range of observational evidence: Studies of classical Cepheids and of the near-infrared luminosity function indicate that about M of stars formed in the GC in the past few tens of Myr (Matsunaga et al. 2011; Nogueras-Lara et al. 2020a). The GC contains three massive ( M) young (2.5-6 Myr) clusters: the Arches and Quintuplet and the central parsec cluster (e.g. Figer et al. 1999; Bartko et al. 2009; Lu et al. 2013; Hosek et al. 2022). There are H II regions and massive YSOs distributed throughout the nuclear disk (e.g. Nandakumar et al. 2018; Hankins et al. 2019). In addition, on the order of 100 apparently isolated massive young stars have been discovered distributed throughout the GC (Dong et al. 2011; Clark et al. 2021). There are about M of young stars in the Sgr B1 H II region (Nogueras-Lara et al. 2022). Sgr B2 is currently the most active site of star formation in the Galaxy. It contains several hundred, highly embedded massive OB stars (e.g., Schmiedeke et al. 2016; Ginsburg et al. 2018).
@@
Figure 4 shows a simulation of a field observed at the GC with F140M, F210M and F480M. The observed CMD becomes severely blurred due to strong and differential extinction, and all stellar populations overlap. Only after de-reddening can we arrive at an interpretable CMD and separate different stellar populations (at least partially). The observed color is a very good tracer of reddening, because the intrinsic colors of all types of stars are small and almost constant with these filters (Fig. 3).
@@
The NSD is the most prolific site of star formation in the Milky Way, but so far we lack a census of young stars in this region. Observational evidence points to the existence of a significant number of clusters and associations with smaller masses or in a more advanced state of dissolution than Arches and Quintuplet ( M, e.g. Dong et al. 2011; Matsunaga et al. 2011; Clark et al. 2021; Nogueras-Lara et al. 2022; Martínez-Arranz et al. 2023). Without NIRCam, they are undetectable, because they do not show up as over-densities in the GC, and it is barely possible to differentiate between massive young stars and older, cool giants in the GC photometrically, as the degeneracy between reddening and the stellar color is significant. While spectroscopy can discern the stellar types, it is impossible to cover the necessary large field with sufficiently sensitive spectroscopic observations at the required angular resolution of .
@@ 3.3 What is the origin and environmental variation of the stellar initial mass function?
Counting the number of YSOs will solve one of the longest-standing questions in Galactic Center star formation research: Why do standard SFR measurements based on IR luminosity, cm-continuum emission and high-mass YSO (HMYSO; 10 M) number counts all show the CMZ SFR is 1–2 orders of magnitude lower than predicted by dense gas star formation relations? As existing SFR measurements are completely dominated by the light from high-mass stars, they miss a population of low-mass YSOs (LMYSOs) left undetected by current facilities. If JWST detects a large population of LMYSOs, the ‘depressed’ CMZ SFR conundrum may be resolved. However, such a solution implies that canonical SFR calibrations, upon which much of our understanding of galaxy evolution is based, underestimate SFRs by over an order of magnitude. If JWST confirms the previous SFR measurements and finds the expected population of LMYSOs extrapolating from known HMYSOs and extended star formation tracers, the SFR calibrations will be validated, but the mystery of GC SFR deepens.
@@
A fundamental prediction of star formation theories combining the scale-free physics of turbulence with gravity is an environmentally dependent volume density threshold for star formation. This ‘critical density’ depends on gas properties such as the mean density, virial ratio, and Mach number (Federrath & Klessen 2012), so is predicted to be 10 times larger in the CMZ than the disk. Combining the JWST census of YSOs with measurements of the physical and kinematic gas properties from existing (sub)millimeter surveys, we will determine how the number and luminosity of YSOs vary with gas column density and velocity dispersion. Measurement of the relation between the YSO surface density and gas properties will determine the value of the environmentally dependent critical density.
@@
It is extremely challenging to distinguish YSOs photometrically from highly reddened main sequence and giant stars. This is an area of open discussion and the community still needs to identify the best strategy to distinguish these populations. Nevertheless, JWST is poised to play a critical role in changing our view of star formation in the Galactic center by enabling star-counting based measurement of the star formation rate and detailed study of how and where stars of different masses form. The current state of the art using Spitzer is limited to the most intrinsically luminous sources (Yusef-Zadeh et al. 2009) and is subject to substantial contamination (Koepferl et al. 2015). Spectroscopic observations have shown that the star formation rates estimated from Spitzer observations are entirely dominated by contaminants (An et al. 2017; Jang et al. 2022), but JWST’s high resolution, especially at 4.8m, will enable us to separate genuine massive YSOs from background sources shining through clouds. JWST should readily detect the vast majority of L YSOs with the planned shallow exposures and is capable of detecting YSOs down to L in many cases (see Appendix A). Further study is needed, however, to determine the most efficient NIRCam filter combinations to distinguish these from deeply embedded post-main-sequence stars.
@@
The 3D distribution of molecular clouds can, on the one hand, be used to interpret the variable signal from X-ray reflecting clouds (e.g. Ponti et al. 2013) and must, on the other hand, be consistent with time variability measurements and polarization measurements (IXPE) in the X-ray domain. The 3D distribution of molecular clouds can be used to see how they constrain and redirect outflows from the innermost regions.
NIRCam astrometry, combined with astrometry from the ground (GALACTICNUCLEUS, see section 5) will provide stellar proper motions with a precision of better than 10 km s (0.25 mas yr). A second epoch of JWST imaging in at least one band will be needed to measure proper motions in the densest parts of clouds, where GALACTICNUCLEUS detects no stars. Combining the JWST proper motions with ALMA radial velocities of clouds, which have relative line-of-sight distances measured, will lead to a determination of the location of the cloud relative to the nucleus as well as provide constraints on the cloud’s 3D space velocity. The identification of young stellar objects (YSOs) embedded in the cloud, combined with the proper motions of these YSOs and the cloud’s radial velocity will lead to a measurement of the cloud’s 3D velocity vector.
@@ 4.2 Filters
The completeness of the proposed observations will be limited by stellar confusion. Since the target is also comparatively bright, narrow and medium-band filters will provide sufficiently high signal-to-noise for all objectives while targeting key diagnostic lines of the ISM and key sections of the stars’ continuum radiation. These filters have the added advantage of well-defined effective wavelengths (which are a function of both stellar type and reddening) and minimised saturation. With the chosen filters we will (hard-)saturate stars brighter than mag. Preliminary work with F212N NIRCam observations of Sgr C have taught us that we can repair the PSFs of at least two magnitude brighter saturated stars. This is both important for obtaining measurements of these stars and for removing the bright PSF features around these stars, which hamper the detection and measurement of fainter stars. Stars brighter than mag make up less than of all stars that we expect to detect in the survey area, assuming the star formation history of Nogueras-Lara et al. (2020a). The surface number density of such stars is arcmin (SIRIUS/IRSF; Nishiyama et al. 2005): The small number and surface density of such bright stars implies that saturation and extended bright star PSFs will have a minimal impact on our science goals.
@@
We propose to carry out the survey with four or five filters: F140M, F187N, F210M, F405N and F480M. Sufficiently deep observations at the shortest wavelength (F140M) may possibly be obtained with HST WFC3 and the F127M filter. Observations with the NIR AND MIR filters will be carried out simultaneously, taking advantage of NIRCams two modules. F210M and F480M are excellent tracers of the stellar continuum. Since the intrinsic color at F210M-F480M of practically all stars brighter than the crowding limit is almost negligible and practically constant (Fig. 4), these filters will allow us to measure the reddening of each star individually. F187N and F405N are centered on the Paschen- and Brackett- H I lines and will allow us to infer the conditions of the ISM (together with the F212N filter, which traces rovibrational emission of H in shocks and photo dominated regions), and find massive post main sequence stars via the line emission in their stellar winds (Dong et al. 2011), and possibly identify YSOs from excess emission from accretion (e.g. Alcalá et al. 2017).

### v7

@@ 2.2 The unique observational capabilities of JWST in Galactic Center context
Angular resolution: Crowding is the main factor that limits the detection of faint stars at the GC, such as old stars near the main sequence turn-off and low-mass YSOs (Fig. 1). JWST provides a ten times higher angular resolution than Spitzer and can deliver almost the resolution of adaptive optics-assisted cameras at ground-based 8 m-class telescopes. However, since precision measurements with the latter are strongly limited by anisoplanatic effects, their useful field-of-view is only approximately 0.25 arcmin2, while the effective field-of-view of NIRCam is roughly 40 times larger in each channel.
@@
Ground-based, adaptive optics-assisted instruments on 10m-class telescopes are no alternative to JWST, because (1) they can cover at most arcminute-wide fields with a systematically changing PSF due to anisoplanatic effects, (2) zero point variations, mostly due to PSF instability, are a problem for multi-epoch observations, and (3) they are about times less sensitive at m than JWST. (1) implies that accurate photometry and astrometry can only be performed over very small fields (0.25 arcmin2), thus making any observations of large fields highly time-consuming. (3) means that less than 10% of the stars detected at m can be detected at m from the ground.
@@ 2.3 JWST will produce transformational measurements
With its high sensitivity JWST can potentially detect YSOs down to Lbol 0.1 in fields where detection completeness is not limited by crowding, such as star-forming regions associated with dark clouds. Current data, mainly from the Spitzer Space Telescope, limit detections to massive YSOs, which are rare and, in the Spitzer data, confused with other sources.
@@
With a minimum of two NIRCam epochs, we can measure the proper motions of an estimated 10 million stars at the GC. No existing or planned mission (including Gaia NIR) is or will be able to achieve this. A five year time base will allow proper motion measurements with an accuracy of at least mas yr-1 (10 km s-1). Here we have considered mag stars and have assumed that alignment between two observing epochs can be done with an accuracy of 1 mas. The proper motions will link the Galactic Center to the rest of the Milky Way where proper motions are measured with Gaia and the Vera Rubin observatory. Proper motions are key to disentangling different stellar structures (bar, nuclear stellar disk, nuclear star cluster), identifying accretion events onto the nuclear star cluster and understanding the structure of the nuclear stellar disk (is it an inner bar?). Even a single epoch JWST survey will already yield about one million proper motions by combination with archival HST (e.g. Dong et al. 2011; Libralato et al. 2021) and ground-based images (from the GALACTICNUCLEUS survey).
@@ 2.4 Survey requirements to address the key open questions
Filters: F140M, F187N, F212N, F405N and F480M to enable accurate dereddening, minimize saturation, trace ISM features, and obtain the best possible color-magnitude diagrams (CMDs). Since confusion is less of an issue at the shortest wavelengths, where extinction is significantly higher, an alternative to F140M imaging with NIRCam may be F127M imaging with WFC3/HST.
@@ 2.5 Synergy with other facilities
ALMA has recently completed the ALMA CMZ Exploration Survey (ACES) in the 3.2 mm continuum and over a dozen spectral lines with 2′′ angular resolution.
@@
The Roman Space Telescope (launch 2027) offers a compelling opportunity to perform a high-cadence survey of the GC region at HST-like resolution (e.g. Terry et al. 2023) for time-domain science and the planne Galactic Plane Survey will image the GC with multiple filters. Because of its worse angular resolution Roman WFI images will hit the crowding limit at brighter magnitudes than JWST NIRCam. The proposed JWST observations will greatly enhance the Roman data by providing star-by-star extinction measurements, much-improved depth in highly crowded and/or extinguished regions, identification of close/blended stars, whose photometry/astrometry will be significantly biased, and an epoch of high-precision astrometry for proper motion measurements.
@@ 3 Solving multiple key science problems with a single data set
The 50050 pc CMZ contains 2 - 7 M⊙ of dense molecular gas, corresponding to % of our Galaxy’s molecular gas. Most of the mass is distributed in a roughly ring-like/elliptical shape of 100 pc () radius around Sgr A*, the so-called 100 pc-stream (Fig. 2, e.g. Henshaw et al. 2022). The CMZ includes exceptional ISM features such as the M⊙ “Brick” (Longmore et al. 2012) and the M⊙ Sgr B2 cloud (de Pree et al. 1995; Schmiedeke et al. 2016; Ginsburg et al. 2018, and references therein). The CMZ occupies of the volume of the Milky Way’s disk and reaches a mean molecular gas density times higher than in the Galactic disk.
The NSD consists of M⊙ of stars in a flat, rotating, disk-like structure that has an effective radius of 100 pc and a scale-height of 40 pc (Launhardt et al. 2002; Schönrich et al. 2015; Henshaw et al. 2022; Sormani et al. 2022). With a mass of M⊙ and an effective radius of about 4 pc, the NSC is the densest concentration of stars in the Galaxy (Schödel et al. 2014b; Feldmeier-Krause et al. 2017; Neumayer et al. 2020). The majority of the stars in the NSD and NSC, , appear to have formed more than 8-10 Gyr ago (Nogueras-Lara et al. 2020a; Schödel et al. 2020) and are the most metal-rich population of old stars in the Galaxy (Schultheis et al. 2021). At the kinematic and density center of the NSC lies the M⊙ super-massive black hole Sagittarius A* (Sgr A*) (Ghez et al. 2008; Gravity Collaboration et al. 2020).
Averaged by volume, the mean star formation rate at the GC is one to two orders of magnitude higher than in the Galactic disk, with extreme conditions that resemble those in high redshift star-forming galaxies (Kruijssen & Longmore 2013). Ongoing and intense past star formation activity is witnessed by a broad range of observational evidence: Studies of classical Cepheids and of the near-infrared luminosity function indicate that about M⊙ of stars formed in the GC in the past few tens of Myr (Matsunaga et al. 2011; Nogueras-Lara et al. 2020a). The GC contains three massive ( M⊙) young (2.5-6 Myr) clusters: the Arches and Quintuplet and the central parsec cluster (e.g. Figer et al. 1999; Bartko et al. 2009; Lu et al. 2013; Hosek et al. 2022). There are H II regions and massive YSOs distributed throughout the nuclear disk (e.g. Nandakumar et al. 2018; Hankins et al. 2019). In addition, on the order of 100 apparently isolated massive young stars have been discovered distributed throughout the GC (Dong et al. 2011; Clark et al. 2021). There are about M⊙ of young stars in the Sgr B1 H II region (Nogueras-Lara et al. 2022). Sgr B2 is currently the most active site of star formation in the Galaxy. It contains several hundred, highly embedded massive OB stars (e.g., Schmiedeke et al. 2016; Ginsburg et al. 2018).
@@
Figure 4 shows a simulation of a field observed at the GC with F140M, F212N and F480M. The observed CMD becomes severely blurred due to strong and differential extinction, and all stellar populations overlap. Only after de-reddening can we arrive at an interpretable CMD and separate different stellar populations (at least partially). The observed color is a very good tracer of reddening, because the intrinsic colors of all types of stars are small and almost constant with these filters (Fig. 3).
@@
The NSD is the most prolific site of star formation in the Milky Way, but so far we lack a census of young stars in this region. Observational evidence points to the existence of a significant number of clusters and associations with smaller masses or in a more advanced state of dissolution than Arches and Quintuplet ( M⊙, e.g. Dong et al. 2011; Matsunaga et al. 2011; Clark et al. 2021; Nogueras-Lara et al. 2022; Martínez-Arranz et al. 2023). Without NIRCam, they are undetectable, because they do not show up as over-densities in the GC, and it is barely possible to differentiate between massive young stars and older, cool giants in the GC photometrically, as the degeneracy between reddening and the stellar color is significant. While spectroscopy can discern the stellar types, it is impossible to cover the necessary large field with sufficiently sensitive spectroscopic observations at the required angular resolution of .
@@ 3.3 What is the origin and environmental variation of the stellar initial mass function?
Counting the number of YSOs will solve one of the longest-standing questions in Galactic Center star formation research: Why do standard SFR measurements based on IR luminosity, cm-continuum emission and high-mass YSO (HMYSO; 10 M⊙) number counts all show the CMZ SFR is 1–2 orders of magnitude lower than predicted by dense gas star formation relations? As existing SFR measurements are completely dominated by the light from high-mass stars, they miss a population of low-mass YSOs (LMYSOs) left undetected by current facilities. If JWST detects a large population of LMYSOs, the ‘depressed’ CMZ SFR conundrum may be resolved. However, such a solution implies that canonical SFR calibrations, upon which much of our understanding of galaxy evolution is based, underestimate SFRs by over an order of magnitude. If JWST confirms the previous SFR measurements and finds the expected population of LMYSOs extrapolating from known HMYSOs and extended star formation tracers, the SFR calibrations will be validated, but the mystery of GC SFR deepens.
@@
A fundamental prediction of star formation theories combining the scale-free physics of turbulence with gravity is an environmentally dependent volume density threshold for star formation. This ‘critical density’ depends on gas properties such as the mean density, virial ratio, and Mach number (Federrath & Klessen 2012), so is predicted to be 103 times larger in the CMZ than the disk. Combining the JWST census of YSOs with measurements of the physical and kinematic gas properties from existing (sub)millimeter surveys, we will determine how the number and luminosity of YSOs vary with gas column density and velocity dispersion. Measurement of the relation between the YSO surface density and gas properties will determine the value of the environmentally dependent critical density.
@@
It is extremely challenging to distinguish YSOs photometrically from highly reddened main sequence and giant stars. This is an area of open discussion and the community still needs to identify the best strategy to distinguish these populations. Nevertheless, JWST is poised to play a critical role in changing our view of star formation in the Galactic center by enabling star-counting based measurement of the star formation rate and detailed study of how and where stars of different masses form. The current state of the art using Spitzer is limited to the most intrinsically luminous sources (Yusef-Zadeh et al. 2009) and is subject to substantial contamination (Koepferl et al. 2015). Spectroscopic observations have shown that the star formation rates estimated from Spitzer observations are entirely dominated by contaminants (An et al. 2017; Jang et al. 2022), but JWST’s high resolution, especially at 4.8m, will enable us to separate genuine massive YSOs from background sources shining through clouds. JWST should readily detect the vast majority of L⊙ YSOs with the planned shallow exposures and is capable of detecting YSOs down to L⊙ in many cases (see Appendix A). Further study is needed, however, to determine the most efficient NIRCam filter combinations to distinguish these from deeply embedded post-main-sequence stars.
@@
The 3D distribution of molecular clouds can, on the one hand, be used to interpret the variable signal from X-ray reflecting clouds (e.g. Ponti et al. 2013) and must, on the other hand, be consistent with time variability measurements and polarization measurements (IXPE) in the X-ray domain (Marin:2023mj; Khabibullin:2025lk). The 3D distribution of molecular clouds can be used to see how they constrain and redirect outflows from the innermost regions.
NIRCam astrometry, combined with astrometry from the ground (GALACTICNUCLEUS, see section 5) will provide stellar proper motions with a precision of better than 10 km s-1 (0.25 mas yr-1). A second epoch of JWST imaging in at least one band will be needed to measure proper motions in the densest parts of clouds, where GALACTICNUCLEUS detects no stars. Combining the JWST proper motions with ALMA radial velocities of clouds, which have relative line-of-sight distances measured, will lead to a determination of the location of the cloud relative to the nucleus as well as provide constraints on the cloud’s 3D space velocity. The identification of young stellar objects (YSOs) embedded in the cloud, combined with the proper motions of these YSOs and the cloud’s radial velocity will lead to a measurement of the cloud’s 3D velocity vector.
@@ 4.2 Filters
The completeness of the proposed observations will be limited by stellar confusion. Since the target is also comparatively bright, narrow and medium-band filters will provide sufficiently high signal-to-noise for all objectives while targeting key diagnostic lines of the ISM and key sections of the stars’ continuum radiation. These filters have the added advantage of well-defined effective wavelengths (which are a function of both stellar type and reddening) and minimised saturation. With the chosen filters we will (hard-)saturate stars brighter than mag. Preliminary work with F212N NIRCam observations of Sgr C have taught us that we can repair the PSFs of at least two magnitude brighter saturated stars. This is both important for obtaining measurements of these stars and, even more so, for removing the bright PSF features around these stars, which hamper the detection and measurement of fainter stars. Stars brighter than mag make up of all stars that we expect to detect in the survey area, assuming the star formation history of Nogueras-Lara et al. (2020a). The small number and surface density of such bright stars implies that saturation and extended bright star PSFs will have a minimal impact on our science goals.
@@
We propose to carry out the survey with four or five filters: F140M, F187N, F212N, F405N and F480M. Sufficiently deep observations at the shortest wavelength (F140M) may possibly be obtained with HST WFC3 and the F127M filter. Observations with the NIR AND MIR filters will be carried out simultaneously, taking advantage of NIRCams two modules. F212N and F480M are excellent tracers of the stellar continuum. Tests with atmosphere models using the SPISEA code (Hosek:2020mi), show that the magnitude difference between the medium band filter F210M and the narrow band filter F212N is mag for the stars that can be detected by NIRCam at the GC. Even though F212N is centred on an H2 line, there is little such emission at the GC, as existing NIRCam observations of the GC show (Hosek M. W. Jr., Ginsburg A.,priv. comm.). Also, PSF fitting tools such as StarFinder will take any diffuse emission underlying point sources into account. Since the intrinsic color at F212N-F480M of practically all stars brighter than the crowding limit is almost negligible and practically constant (Fig. 4), these filters will allow us to measure the reddening of each star individually. F187N and F405N are centered on the Paschen- and Brackett- H I lines and will allow us to infer the conditions of the ISM (together with the F212N filter, which traces rovibrational emission of H2 in shocks and photo dominated regions), and find massive post main sequence stars via the line emission in their stellar winds (Dong et al. 2011), and possibly identify YSOs from excess emission from accretion (e.g. Alcalá et al. 2017).